4.If you purchase a house today for $175,000, what will be your mortgage payment if you paid 15% down and borrowed the balance for 30 years at 9% compounded monthly?

  • A) $1,408
  • B) $1,419
  • C) $1,207
  • D) $1,197
Answer #1
  • D) $1,197
Monthly payment = Loan amount / Present value of annuity of 1
= $       1,48,750 / 124.2819
= $             1,197
Working;
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= 124.2818657 i = 9%/12 = 0.0075
n = 30*12 = 360
Loan amount = Cost of house - Down Payment
= $       1,75,000 - (175000*15%)
= $       1,75,000 - $ 26,250
= $       1,48,750